evaluate the following integral
∫ sec x tan x / 3 sec x + 5 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{secxtanx}{3secx+5}dx
3secx+5 = t
3secxtanxdx = dt
I = \frac{1}{3}\int \frac{1}{t}dt
I = \frac{1}{3}lnt + c
I = \frac{1}{3}ln(3secx+5) + c